the sum of three consecutive even integers is 36. Find the smallest of them.
Answer by nerdybill(7384)

the sum of three consecutive even integers is 36. Find the smallest of them.
.
Let x = smallest of three consecutive even integers
then
x+2 = second integer
x+4 = third
.
x + x+2 + x+4 = 36
3x + 6 = 36
3x = 30
x = 10 (answer)